3.a. Two goods have a cross-price elasticity of demand of +1.2 (a) would you describe the 

goods as substitutes or complements? (b) If the price of one of the goods rises by 5 per 

cent, what will happen to the demand for the other good, holding other factors constant?

(5 Marks)

3.b. Calculate Marginal Utility and Average Utility from the information given in the below 

table: (5 Marks)

Quantity Consumed Total Utility

1 20

2 35

3 47

4 55

5 60


1
Expert's answer
2021-10-24T20:15:25-0400

(a) The goods are substitute goods since they have a positive cross-price elasticity of demand.


(b) If the price of one of the two goods increases by 5 per cent, it will be substi­tuted by the other good so that the quantity demanded of this other good will increase.

With positive cross-price elasticity being equal to 1.2, the quantity demanded of the other good will increase by "1.2\times5\%=6\%" = 6 per cent


(c)Marginal utility and average utility

Marginal Utility= change in total utility divided by change in quantity consumed


Average Utility= total utility divided by total quantity consumed


Quantity Consumed Total Utility Marginal utility Average utility

1 20 ="\frac{20}{1}" =20 "=\frac{20}{1}=20"


2 35 ="\frac{35-20}{2-1}" =15 "=\frac{35}{2}=17.5"


3 47 ="\frac{47-35}{3-2}" =12 "=\frac{47}{3}=15.67"

4 55 ="\frac{55-47}{4-1}" =8 "=\frac{55}{4}=13.75"


5 60 ="\frac{60-55}{5-4}" =5 "=\frac{60}{5}=12"